# parskip=half

### v2

@@ Abstract
Causal survival analysis combines survival analysis and causal inference to evaluate the effect of a treatment or intervention on a time-to-event outcome, such as survival time. It offers an alternative to relying solely on Cox models for assessing these effects. In this paper, we present a comprehensive review of estimators for the average treatment effect measured with the restricted mean survival time, including regression-based methods, weighting approaches, and hybrid techniques. We investigate their theoretical properties and compare their performance through extensive numerical experiments. Our analysis focuses on the finite-sample behavior of these estimators, the influence of nuisance parameter selection, and their robustness and stability under model misspecification. By bridging theoretical insights with practical evaluation, we aim to equip practitioners with both state-of-the-art implementations of these methods and practical guidelines for selecting appropriate estimators for treatment effect estimation. Among the approaches considered, G-formula two-learners, AIPCW-AIPTW, Buckley-James estimators, and causal survival forests emerge as particularly promising.
@@ 1.1 Context and motivations
Causal survival analysis is a growing field that integrates causal inference (Rubin,, 1974; Hernán MA,, 2020) with survival analysis (Kalbfleisch and Prentice,, 2002) to evaluate the impact of treatments on time-to-event outcomes, while accounting for censoring situations where only partial information about an event’s occurrence is available.
Although it is a relatively new domain, the literature has quickly expanded—often in a fragmented way. Researchers face a wide array of estimators and must make several methodological decisions, often without clear guidance. Moreover, many of the proposed methods lack accessible implementations. This is particularly important given that, in many real-world scenarios where randomized trials are infeasible, well-conducted observational studies—with explicit and plausible causal assumptions—remain the only viable source of evidence about the treatment effect (Makary and Prasad,, 2025). This calls for a structured review that consolidates available approaches, clarifies their assumptions, and provides an evaluation of estimator performance to make informed methodological choices — particularly in finite sample settings. Such a survey also has the potential to help identify remaining methodological challenges that need to be addressed.
@@
Related efforts have already been made to clarify methodological options in survival analysis. For instance, Denz et al., (2023) focus on confounders-adjusted survival curves under proportional hazards and explore model misspecification in that context. However, existing reviews often remain within the Cox model framework, and may not fully address the diversity of estimators and causal estimands available—particularly those applicable when the proportional hazards assumption does not hold or when censoring is not independent.
@@
Despite its widespread use in biomedical research, we decided not to cover the estimation of the Hazard Ratio (HR) which presents several well-documented limitations. First, when using Cox model, it relies on the proportional hazards assumption, which is not necessarely satisfied in practice. Second, hazard ratios are generally not causally interpretable—even in randomized controlled trials—due to built-in selection bias that can distort the comparison between groups unless the treatment has no effect at all (Martinussen and Vansteelandt,, 2013; Martinussen et al.,, 2020). Third, the HR is a non-collapsible measure, meaning that the effect estimated in the entire population (the marginal effect) can differ from the effect estimated within subgroups (the conditional effect), even when there is no confounding (Huitfeldt et al.,, 2019; Greenland et al.,, 1999). Finally, the HR could be difficult to interpret clinically, particularly because it reflects an instantaneous risk rather than a cumulative or absolute survival benefit (Averbuch et al.,, 2025). In contrast, survival curves are more informative than HRs and can be easily generated (Hernán,, 2010).
For all these reasons, we focus in this paper on the estimation of the Restricted Mean Survival Time (RMST), a popular causal measure in survival analysis that offers an intuitive interpretation of the average survival time over a specified time horizon.
Numerous methods have been developed for causal survival analysis, including approaches based on pseudo-observations (Andersen et al.,, 2017; Wang,, 2018; van Hage et al.,, 2024) and targeted maximum likelihood estimation (TMLE) (Stitelman et al.,, 2012; Talbot et al.,, 2025). In this work, we focus on estimators based on outcome regression, inverse probability weighting, and hybrid strategies.
1.2 Definition of the estimand: the RMST
@@
We set the analysis in the potential outcome framework, where a patient, described by a vector of covariates $X\in\mathbb{R}^{p}$, either receives a treatment ($A=1$) or is in the control group ($A=0$). The patient will then survive up to a certain time $T^{(0)}\in\mathbb{R}^{+}$ in the control group, or up to a time $T^{(1)}\in\mathbb{R}^{+}$ in the treatment group. In practice, we cannot simultaneously have access to $T^{(0)}$ and $T^{(1)}$, as one patient is either treated or control, but only to $T$ defined as follows:
@@
(Stable Unit Treatment Value Assumption — SUTVA)
$T=AT^{(1)}+(1-A)T^{(0)}$
Due to potential censoring, the outcome $T$ is not completely observed. The most common form of censoring is right-censoring (also known as type II censoring), which occurs when the event of interest has not taken place by the end of the observation period, indicating that it may have occurred later if the observation had continued (Turkson et al.,, 2021). We focus in this study on this type of censoring only and we assume that we observe $\tilde{T}=T\wedge C=\min(T,C)$ for some censoring time $C\in\mathbb{R}^{+}$. When an observation is censored, the observed time is equal to the censoring time.
We also assume that we know whether an outcome is censored or not. In other words, we observe the event status variable $\Delta=\mathbb{I}\{T\leqslant C\}$, where $\mathbb{I}\{\cdot\}$ is the indicator function. $\Delta$ is $1$ if the true outcome is observed, and 0 if it is censored.
We assume observing a $n$-sample of variables $(X,A,\widetilde{T},\Delta)$ stemming from an $n$-sample of the partially unobservable variables $(X,A,T^{(0)},T^{(1)},C)$. A toy example of such data is given in Table 1.
Table 1: Example of a survival dataset. Each observation corresponds to a single patient. In practice, only $X$, AA, $\Delta$, and $\widetilde{T}$ are observed. The column Obs. Outcome represents the observed survival outcome.
@@
Our aim is to estimate the Average Treatment Effect (ATE) defined as the difference between the Restricted Mean Survival Time (Royston and Parmar 2013).
@@
(Causal effect: Difference between Restricted Mean Survival Time)
$\theta_{\mathrm{RMST}}=\mathbb{E}\left[T^{(1)}\wedge\tau-T^{(0)}\wedge\tau\right],$
where $a\wedge b:=\min(a,b)$ for $a,b\in\mathbb{R}$, and $\tau>0$ is a pre-specified time horizon over which survival is evaluated.
@@
Consequently, $\theta_{\mathrm{RMST}}$ can be interpreted as the mean difference between the survival function of treated and control until a fixed time horizon $\tau$. A RMST difference $\theta_{\mathrm{RMST}}=10$ days with $\tau=200$ means that, on average, the treatment increases the survival time by 10 days at 200 days. We give a visual interpretation of RMST in Figure 1.
Figure 1: Plot of the estimated survival curves on synthetic toy-data. The $\theta_{\mathrm{RMST}}$ at $\tau=50$ corresponds to the yellow shaded area between the two survival curves. The curves have been estimated using the Kaplan-Meier estimator, see Section 2.1.
Although the present work focuses on the estimation of the difference in RMST, we would like to stress that the causal effect can be assessed through other measures, such as for instance the difference between the survival functions
$\theta_{\mathrm{SP}}:=S^{(1)}(\tau)-S^{(0)}(\tau)$
for some time $\tau$, see for instance (Ozenne et al.,, 2020). As mentioned in Section 1.1, another widely used measure (though not necessarily causal) is the hazards ratio, defined as
$\theta_{\mathrm{HR}}:=\frac{\lambda^{(1)}(\tau)}{\lambda^{(0)}(\tau)},$
where the hazard function $\lambda^{(a)}$ is defined as
$\lambda^{(a)}(t):=\lim_{h\to 0^{+}}\frac{\mathbb{P}(T^{(a)}\in[t,t+h]|T^{(a)}\geqslant t)}{h}.$
in a continuous setting, or as $\lambda^{(a)}(t):=\mathbb{P}(T^{(a)}=t|T^{(a)}\geqslant t)$ when the survival times are discrete. The hazard functions and the survival functions are linked through the identities
$S^{(a)}(t)=\exp\left(-\Lambda^{(a)}(t)\right)\quad\text{where}\quad\Lambda^{(a)}(t):=\int_{0}^{t}\lambda^{(a)}(s)\mathop{}\!\mathrm{d}s,$} (2)
in the continuous case. The functions $\Lambda^{(a)}$ are called the cumulative hazard functions. In the discrete case, we have in turn
$S^{(a)}(t)=\prod_{t_{k}\leqslant t}\left(1-\lambda^{(a)}(t_{k})\right),$} (3)
where $\{t_{1},\dots,t_{K}\}$ are the discrete event times of $T^{(a)}$. These hazard functions are classically used to model the survival times and the censoring times, see Section 2.2.1.
1.3 Organization of the paper
In this paper, we detail the minimal theoretical framework that allows the analysis of established RMST estimators in the context of both Randomized Controlled Trials (Section 2) and observational data (Section 3). We give their statistical properties (consistency, asymptotic normality) along with proofs when possible. We then conduct in Section 6 a numerical study of these estimators through simulations under various experimental conditions, including independent and conditionally independent censoring and correct and incorrect model specifications. We conclude in Section 7 with practical recommendations on estimator selection, based on criteria such as asymptotic behavior, computational complexity, and efficiency.
1.4 Notations
We provide in Table LABEL:tbl-reminder_notations a summary of the notation used throughout the paper.
@@ 1.4 Notations
$\Lambda^{(a)},a\in\{0,1\}$ Cumulative hazard function of the potential survival times
@@
$\widetilde{T}$ Observed time ($T\wedge C$)
@@ 2 Causal survival analysis in Randomized Controlled Trials
Randomized Controlled Trials (RCTs) are the gold standard for establishing the effect of a treatment on an outcome, because treatment allocation is controlled through randomization, which ensures (asymptotically) the balance of covariates between treated and controls, and thus avoids problems of confounding between treatment groups. The core assumption in a classical RCT is the random assignment of the treatment (D. B. Rubin 1974).
@@
(Random Treatment Assignment)
$A\perp\mkern-9.5mu\perp T^{(0)},T^{(1)},X$
We also assume that there is a positive probability of receiving the treatment, which we rephrase under the following assumption.
@@
(Trial Positivity)
$0<\mathbb{P}(A=1)<1$
Under Assumptions 1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) and 2.2 (.Causal survival analysis in Randomized Controlled Trials), classical causal identifiability equations can be written to express $\theta_{\mathrm{RMST}}$ without potential outcomes.
@@
However, Equation 4 still depends on $T$, which remains only partially observed due to censoring. To ensure that censoring does not compromise the identifiability of treatment effects, we must impose certain assumptions on the censoring process, standards in survival analysis, namely, independent censoring and conditionally independent censoring. These assumptions lead to different estimation approaches. We focus on two strategies: those that aim to estimate $\mathbb{E}[T\wedge\tau|A=a]$ directly (e.g., through censoring-unbiased transformations, see Section 2.2), and those that first estimate the survival curves to derive RMST via Equation 1 (such as the Kaplan-Meier estimator and all its variants, see the next Section).
@@ 2.1 Independent censoring: the Kaplan-Meier estimator
In a first approach, one might assume that the censoring times are independent on the rest of the variables.
@@
(Independent Censoring)
$C\perp\mkern-9.5mu\perp T^{(0)},T^{(1)},X,A.$
Under Assumption 2.3, subjects censored at time $t$ are representative of all subjects who remain at risk at time $t$.
Figure 2 represents the causal graph when the study is randomized and outcomes are observed under independent censoring. The treatment assignment $A$ is randomized and hence independent of baseline covariates $X$, which may influence the potential survival outcomes $T^{(a)}$. The observed survival time $T$ depends on both $A$ and $X$, while the censoring time $C$ is assumed to be conditionally independent of the potential outcomes given $X$, reflecting an independent censoring mechanism. This structure encodes key assumptions used throughout this section, including consistency, random treatment assignment, and independent censoring.
@@
We also assume that there is no almost-sure upper bound on the censoring time before $\tau$, which we rephrase under the following assumption.
@@
(Positivity of the Censoring Process)
There exists $\varepsilon>0$ such that
$G(t)\geqslant\varepsilon\quad\text{for all}\quad t\in[0,\tau).$
If indeed it was the case that $\mathbb{P}(C<t)=1$ for some $t<\tau$, then we would not be able to infer anything on the survival function on the interval $t,\tau]$ as all observation times $\widetilde{T}_{i}$ would be in $[0,t]$ almost surely. In practice, adjusting the threshold time $\tau$ can help satisfy the positivity assumption. For instance, in a clinical study, if a subgroup of patients has zero probability of remaining uncensored at a given time, $\tau$ can be modified to ensure that participants have a feasible chance of remaining uncensored up to the revised threshold.
The two Assumptions 2.3 (.Independent censoring: the Kaplan-Meier estimator) and 2.4 (.Independent censoring: the Kaplan-Meier estimator) together allow the distributions of $T^{(a)}$ to be identifiable, in the sense that there exists an identity that expresses $S^{(a)}$ as a function of the joint distribution of $(\widetilde{T},\Delta,A=a)$, see for instance (Ebrahimi et al.,, 2003) for such a formula in a non-causal framework. This enables several estimation strategies, the most well-known being the Kaplan-Meier product-limit estimator.
To motivate the definition of the latter and make explicit the identifiability identity, we set the analysis in the discrete case. We let $\{t_{k}\}_{k\geqslant 1}$ be a set of positive and increasing times and assume that $T\in\{t_{k}\}_{k\geqslant 1}$ almost surely. Then for any $t\in[0,\tau]$, it holds, letting $t_{0}=0$ by convention, thanks to Equation 3,
$S(t|A=a)$ $=\mathbb{P}(T>t|A=a)=\prod_{t_{k}\leqslant t}\left(1-\mathbb{P}(T=t_{k}|T>t_{k-1},A=a)\right)$
$=\prod_{t_{k}\leqslant t}\left(1-\frac{\mathbb{P}(T=t_{k},A=a)}{\mathbb{P}(T\geqslant t_{k},A=a)}\right).$
Using Assumptions 2.3 (.Independent censoring: the Kaplan-Meier estimator) and 2.4 (.Independent censoring: the Kaplan-Meier estimator), we find that
$\frac{\mathbb{P}(T=t_{k},A=a)}{\mathbb{P}(T\geqslant t_{k},A=a)}=\frac{\mathbb{P}(T=t_{k},C\geqslant t_{k},A=a)}{\mathbb{P}(T\geqslant t_{k},C\geqslant t_{k},A=a)}=\frac{\mathbb{P}(\widetilde{T}=t_{k},\Delta=1,A=a)}{\mathbb{P}(\widetilde{T}\geqslant t_{k},A=a)},$} (5)
yielding the final identity
$S(t|A=a)=\prod_{t_{k}\leqslant t}\left(1-\frac{\mathbb{P}(\widetilde{T}=t_{k},\Delta=1,A=a)}{\mathbb{P}(\widetilde{T}\geqslant t_{k},A=a)}\right).$} (6)
Notice that the right hand side only depends on the distribution of the observed tuple $(A,\widetilde{T},\Delta)$. This last equation suggests in turn to introduce the quantities
$D_{k}(a):=\sum_{i=1}^{n}\mathbb{I}(\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A=a)\quad\text{and}\quad N_{k}(a):=\sum_{i=1}^{n}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A=a),$} (7)
which correspond respectively to the number of deaths $D_{k}(a)$ and of individuals at risk $N_{k}(a)$ at time $t_{k}$ in the treated group (a=1) or in the control group (a=0).
@@
(Kaplan-Meier estimator, (Kaplan and Meier,, 1958)) With $D_{k}(a)$ and $N_{k}(a)$ defined in Equation 7, we let
$\widehat{S}_{\mathrm{KM}}(t|A=a):=\prod_{t_{k}\leqslant t}\left(1-\frac{D_{k}(a)}{N_{k}(a)}\right).$} (8)
The associated RMST estimator is then simply defined as
$\widehat{\theta}_{\mathrm{KM}}=\int_{0}^{\tau}\widehat{S}_{KM}(t|A=1)-\widehat{S}_{KM}(t|A=0)dt.$} (9)
The Kaplan-Meier estimator is the Maximum Likelihood Estimator (MLE) of the survival functions, see for instance Kaplan and Meier, (1958). Furthermore, because $D_{k}(a)$ and $N_{k}(a)$ are sums of i.i.d. random variables, the Kaplan-Meier estimator inherits some convenient statistical properties. In Annex 9 (Section 9.1) we show that it exhibits exponentially fast bias decay (Proposition 9.1), asymptotic normality with explicitly characterized variances (Proposition 9.2). We finally mention that the KM estimator as defined in Definition 2.1 still makes sense in a non-discrete setting, and one only needs to replace the fixed grid $\{t_{k}\}$ by the values at which we observed an event $(\widetilde{T}_{i}=t_{k},\Delta_{i}=1)$. We refer to Breslow, (1974) for a study of this estimator in the continuous case and to Aalen et al., (2008), Sec 3.2 for a general study of the KM estimator through the prism of point processes.

### v3

@@ Abstract
The restricted mean survival time (RMST) difference offers an interpretable causal contrast to estimate the treatment effect for time-to-event outcomes, yet a wide range of available estimators leaves limited guidance for practice. We provide a unified review of RMST estimators for randomized trials and observational studies, establish identification and asymptotic properties, and supply new derivations where needed. Our extensive simulation study compares simple nonparametric methods (such as unweighted Kaplan–Meier estimators) alongside parametric and nonparametric implementations of the G-formula, weighting approaches, Buckley–James transformations, and augmented estimators under diverse censoring mechanisms and model specifications. Across scenarios, classical Kaplan–Meier estimators (weighted when required by the censoring process) and G-formula methods perform well in randomized settings, while in observational data G-formula estimators remain competitive; however, augmented estimators such as AIPTW–AIPCW generally offer robustness to model misspecification and a favorable bias–variance trade-off. Parametric estimators perform best under correct specification, whereas nonparametric methods avoid functional assumptions but require large sample sizes to achieve reliable performance. We offer practical recommendations for estimator choice and provide open-source R code to support reproducibility and application.
@@ 1 Introduction
Causal survival analysis is a rapidly evolving field that integrates causal inference (rubin_1974; Hernan2020) with survival analysis (kalbfleisch2002) to evaluate the impact of treatments on time-to-event outcomes, while accounting for censoring, i.e. situations where only partial information about an event’s occurrence is available. In this litterature, the Restricted Mean Survival Time (RMST) difference has emerged as a promising causal measure that offers an intuitive interpretation of the average survival time over a specified time horizon.
Despite its recent emergence, the field has expanded quickly, though often in a fragmented manner. As a result, researchers are faced with a broad spectrum of estimators, from weighting and regression-based approaches to more complex hybrid strategies, and must navigate numerous methodological decisions. These include, for example, the choice of how to estimate nuisance components, whether through parametric models, or nonparametric techniques, frequently without clear guidance. The situation is further complicated by the limited availability of accessible software implementations. Addressing these challenges is especially crucial in real-world contexts where randomized trials are not feasible and carefully designed observational studies, grounded in explicit causal assumptions, represent the primary source of evidence on treatment effects (Makary2025). These considerations underscore the need for a structured review that consolidates existing approaches, clarifies their underlying assumptions, and evaluates estimator performance to support informed methodological decisions, particularly in finite-sample settings. Such a survey could also help to identify key methodological gaps that warrant further research.
@@
Related efforts have already been made to clarify methodological options in survival analysis. For instance, denz2023 focus on confounders-adjusted survival curves under proportional hazards and assess the impact of model misspecification. Yet, most existing reviews remain confined to the Cox model framework and thus overlook the broader range of estimators and causal estimands now available.
@@
Despite its widespread use, we chose not to cover the estimation of the hazard ratio (HR), typically derived from a Cox model, because of several well-documented limitations. It relies on the proportional hazards assumption, which may not hold in practice. Moreover, the HR does not represent a causal contrast (Martinussen2020), even in randomized trials, due to inherent selection bias that can distort group comparisons unless the treatment has no effect (Martinussen2013; Hernan2010_HR). The HR is also non-collapsible, meaning marginal effects cannot be inferred from conditional ones (huitfeldt2019collapsibility; greenland1999). Finally, its clinical interpretation could be difficult, as it reflects an instantaneous risk rather than a cumulative or absolute survival benefit (averbuch2025).
In this paper, after introducing the notations and the definition of the restricted mean survival time (Section 1.1), we present the statistical framework required to define and estimate RMST in both randomized controlled trials (Section 2) and observational settings (Section 3). We characterize their key statistical properties, such as consistency and asymptotic normality, and provide a unified, self-contained set of proofs. Where existing results are incomplete or unavailable, we supply original derivations; where they exist, we rederive them within a common formalism that enhances clarity and facilitates direct comparison between estimators. In Section 5, we complement this theoretical analysis with a simulation study assessing estimator performance under various conditions, including independent and conditionally independent censoring and both correctly specified and misspecified models. Finally, in Section 6, we offer practical recommendations for estimator selection, grounded in considerations such as finite-sample behavior and computational complexity.
1.1 Notations and definition of the Restricted Mean Survival Time
@@
We set the analysis in the potential outcome framework, where a patient, described by a vector of covariates $X\in\mathbb{R}^{p}$, either receives a treatment ($A=1$) or is in the control group ($A=0$). The patient’s survival time would then be $T^{(0)}\in\mathbb{R}^{+}$ under control and $T^{(1)}\in\mathbb{R}^{+}$ under treatment. In practice, both potential outcomes cannot be observed simultaneously for the same individual, since each patient receives only one treatment condition. Instead, we observe the realized survival time $T$, defined as follows:
@@
(Stable Unit Treatment Value Assumption — SUTVA) $T=AT^{(1)}+(1-A)T^{(0)}$.
Because of potential censoring, the outcome $T$ is not always fully observed. The most common form is right-censoring, which occurs when the event has not occurred by the end of follow-up (Turkson2021). In what follows, we focus on this case: for each subject, we observe an $n$-sample $(X,A,\tilde{T},\Delta)$, where $\tilde{T}=\min(T,C)$, $C\in\mathbb{R}_{+}$, and $\Delta=\mathbb{I}\{T\leq C\}$ (equal to $1$ if the event time is observed and 0 otherwise). These observed variables are derived from the unobservable tuple $(X,A,T^{(0)},T^{(1)},C)$ illustrated in Table 1.
Table 1: Example of a survival dataset. Each observation corresponds to a patient. In practice, only $X$, AA, $\Delta$, and $\widetilde{T}$ are observed. The column Obs. Outcome represents the observed survival outcome.
@@
Our aim is to estimate the Average Treatment Effect (ATE) defined as the difference between the Restricted Mean Survival Time (royston2013).
@@
(Causal effect: Difference between Restricted Mean Survival Time) We define $\theta_{\mathrm{RMST}}=\mathbb{E}\!\left[T^{(1)}\wedge\tau-T^{(0)}\wedge\tau\right],$ where $a\wedge b:=\min(a,b)$ for $a,b\in\mathbb{R}$, and $\tau>0$ is a pre-specified time horizon over which survival is evaluated.
@@
Consequently, $\theta_{\mathrm{RMST}}$ can be interpreted as the mean difference between the survival function of treated and control until a fixed time horizon $\tau$ (See Figure 1). A RMST difference $\theta_{\mathrm{RMST}}=10$ days with $\tau=200$ means that, on average, the treatment increases the survival time by 10 days at 200 days.
Although the present work focuses on the estimation of the difference of RMST, we would like to stress that the causal effect can be assessed through other measures, such as for instance the difference between the survival functions $\theta_{\mathrm{SP}}:=S^{(1)}(\tau)-S^{(0)}(\tau)$ for some time $\tau$ (Ozenne20_AIPTW_AIPCW).
Figure 1: Estimated survival curves on synthetic data. The $\theta_{\mathrm{RMST}}$ at $\tau=50$ corresponds to the yellow area between them. Curves are estimated with the Kaplan-Meier estimator, see Section 2.1.
We provide in Table LABEL:tbl-reminder_notations a summary of the notations used throughout the paper.
@@
$\Lambda^{(a)},a\in\{0,1\}$ Cumulative hazard function $\Lambda^{(a)}(t):=\int_{0}^{t}\lambda^{(a)}(s)\mathop{}\!\mathrm{d}s$ of the potential survival times
@@
$\widetilde{T}$ Observed time ($T\wedge C=\min(T,C)$)
@@ 2 Causal survival analysis in Randomized Controlled Trials
Randomized Controlled Trials (RCTs) are the gold standard for establishing the effect of a treatment on an outcome, because treatment allocation is controlled through randomization, which ensures (asymptotically) the balance of covariates between treated and controls, and thus avoids problems of confounding between treatment groups. In a classical RCT, treatment is randomly assigned (rubin_1974), and everyone could, in principle, be assigned to either arm.
@@
(Random Treatment Assignment) $A\perp\mkern-9.5mu\perp T^{(0)},T^{(1)},X$.
@@
(Trial Positivity) $0<\mathbb{P}(A=1)<1$.
Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials) and 2.2 (.Causal survival analysis in Randomized Controlled Trials), classical causal identifiability equations can be written to express $\theta_{\mathrm{RMST}}$ without potential outcomes.
@@
However, Equation 2 still depends on $T$, which is only partially observed due to censoring. To ensure identifiability of treatment effects, standard assumptions on the censoring mechanism are required, namely independent and conditionally independent censoring.
When censoring is independent, standard survival estimators, such as the Kaplan–Meier estimator, directly accommodate censoring, allowing RMST to be obtained from the estimated survival curve via Equation 1 (Section 2.1). Under the more general condition of conditionally independent censoring, censoring-unbiased transformations of the event time can be used to recover $\mathbb{E}[T\wedge\tau\mid A=a]$ (Section 2.2). These transformations enable two estimation strategies: (i) direct estimation, by computing the transformed outcome and averaging it and (ii) adjusted survival-curve estimation, in which the transformation is incorporated within a survival estimator, and RMST is then derived using Equation 1. These two families of methods form the basis of the estimators reviewed in the next sections.
@@ 2.1 Independent censoring: the Kaplan-Meier estimator
(Independent Censoring) $C\perp\mkern-9.5mu\perp T^{(0)},T^{(1)},X,A.$
Under Assumption 2.3, subjects censored at time $t$ are representative of all subjects who remain at risk at time $t$. Figure 2 shows the causal graph where the treatment assignment is random, the time $T$ depends on $(A,X)$, and the censoring time $C$ is independent of all variables.
@@
We assume that there is no almost-sure upper bound on the censoring time before $\tau$.
@@
(Positivity of the Censoring Process) There exists $\varepsilon>0$ such that $G(\tau)\geqslant\varepsilon$.
If indeed it was the case that $\mathbb{P}(C<t)=1$ for some $t<\tau$, then we would not be able to infer anything on the survival function on the interval $t,\tau]$ as all observation times $\widetilde{T}_{i}$ would be in $[0,t]$ almost surely. In practice, adjusting the threshold time $\tau$ can help satisfy this positivity assumption.
Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials), 2.2 (.Causal survival analysis in Randomized Controlled Trials), 2.3 (.Independent censoring: the Kaplan-Meier estimator), and 2.4 (.Independent censoring: the Kaplan-Meier estimator), the potential-outcome survival $S^{(a)}(t)$ is identified as a functional of the joint law of ($\tilde{T},\Delta,A$). This enables several estimation strategies, the most well-known being the Kaplan-Meier, the Maximum Likelihood Estimator (MLE) of the survival functions, see for instance Kaplan1958_1.
@@
(Kaplan-Meier estimator, (Kaplan1958_1))
$\widehat{S}_{\mathrm{KM}}(t|A=a):=\prod_{t_{k}\leqslant t}\left(1-\frac{D_{k}(a)}{N_{k}(a)}\right).$
with $D_{k}(a):=\sum_{i=1}^{n}\mathbb{I}(\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A=a)\quad\text{and}\quad N_{k}(a):=\sum_{i=1}^{n}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A=a),$ which correspond respectively to the number of deaths $D_{k}(a)$ and of individuals at risk $N_{k}(a)$ at time $t_{k}$ in the treated group ($a=1$) or in the control group ($a=0$).
Furthermore, since $D_{k}(a)$ and $N_{k}(a)$ are sums of i.i.d. random variables, the Kaplan–Meier estimator inherits convenient statistical properties. In Appendix A.1.1. Proofs of Section 2.1, we show that it exhibits exponentially fast bias decay (Proposition 6.1) and asymptotic normality with explicitly characterized variances (Proposition 6.2). Finally, the estimator in Definition 2.1 extends naturally to the continuous setting by replacing the fixed grid $\{t_{k}\}$ with observed event times $(\widetilde{T}_{i}=t_{k},\Delta_{i}=1)$. We refer to Breslow1974 for a study of this version and to (Aalen2008, Sec. 3.2) for a broader point-process perspective.
The associated RMST estimator is then defined as
$\widehat{\theta}_{\mathrm{KM}}=\int_{0}^{\tau}\widehat{S}_{\mathrm{KM}}(t|A=1)-\widehat{S}_{\mathrm{KM}}(t|A=0)dt.$} (3)
This estimator inherits KM properties. It is consistent for the target RMST and its bias decays exponentially fast. In RCTs, the independence of arms delivers asymptotic normality with variance equal to the sum of the arm-specific RMST variances.